$SPAD/src/input efi.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool efi.output
)set message test on
)set message auto off
)clear all
--S 1 of 15
EFI:=Expression Integer
--R
--R
--R   (1)  Expression(Integer)
--R                                                                 Type: Domain
--E 1
--S 2 of 15
ber:=operator 'ber
--R
--R
--R   (2)  ber
--R                                                          Type: BasicOperator
--E 2
--S 3 of 15
s:=operator 's
--R
--R
--R   (3)  s
--R                                                          Type: BasicOperator
--E 3
-- s:OP EFI:=operator 's
--S 4 of 15
br:LIST EFI->EFI
--R
--R                                                                   Type: Void
--E 4

———-

 br(x|(x.1)~=0)==(x.1)/(exp((x.1))-1)
 br has a removable singularity at the origin
 br(x|(x.1)=0)== limit(br([y]),y=0)

— * —

--S 5 of 15
br(x) ==
 (x.1) = 0 => limit(br([y]),y=0)
 (x.1)/(exp((x.1))-1)
--R
--R                                                                   Type: Void
--E 5
--S 6 of 15
br([1])
--R
--R   Compiling function br with type List(Expression(Integer)) ->
--R      Expression(Integer)
--R
--R           1
--R   (6)  ------
--R        %e - 1
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 15
br([0])
--R
--R
--R   (7)  1
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 15
fJ:List FRAC INT -> EFI
--R
--R                                                                   Type: Void
--E 8
--S 9 of 15
J(i:PI,j:PI):EFI==ber(s(i)-s(j))
--R
--R   Function declaration J : (PositiveInteger,PositiveInteger) ->
--R      Expression(Integer) has been added to workspace.
--R                                                                   Type: Void
--E 9

--S 10 of 15
function(J(1,2),'fJ,['s])
--R
--R   Compiling function J with type (PositiveInteger,PositiveInteger) ->
--R      Expression(Integer)
--R
--R   (10)  fJ
--R                                                                 Type: Symbol
--E 10
--S 11 of 15
evaluate(ber,br)$BOP1(EFI);
--R
--R
--R                                                          Type: BasicOperator
--E 11
--S 12 of 15
ss:=[1,2]
--R
--R
--R   (12)  [1,2]
--R                                                  Type: List(PositiveInteger)
--E 12
--S 13 of 15
fJ(ss)
--R
--R   Compiling function fJ with type List(Fraction(Integer)) ->
--R      Expression(Integer)
--R
--R           %e
--R   (13)  ------
--R         %e - 1
--R                                                    Type: Expression(Integer)
--E 13
--S 14 of 15
ss:=[1,1]
--R
--R
--R   (14)  [1,1]
--R                                                  Type: List(PositiveInteger)
--E 14
-- fJ doesn't know about the special definition at the origin
--S 15 of 15
fJ(ss)
--R
--R

--R   (15)  1
--R                                                    Type: Expression(Integer)
--E 15
)spool
)lisp (bye)

———-

References^{[1] nothing}

7